# Revisiting LLM Evaluation through Mechanism Interpretability: a New Metric and Model Utility Law

### v1

@@ Abstract
Large Language Models (LLMs) have become indispensable across academia, industry, and daily applications, yet current evaluation methods struggle to keep pace with their rapid development. In this paper, we analyze the core limitations of traditional evaluation pipelines and propose a novel metric, the Model Utilization Index (MUI), which introduces mechanism interpretability techniques to complement traditional performance metrics. MUI quantifies the extent to which a model leverages its capabilities to complete tasks. The core idea is that to assess an LLM’s overall ability, we must evaluate not only its task performance but also the effort expended to achieve the outcome. Our extensive experiments reveal an inverse relationship between MUI and performance, from which we deduce a common trend observed in popular LLMs, which we term the “Utility Law." Based on this, we derive four corollaries that address key challenges, including training judgement, the issue of data contamination, fairness in model comparison, and data diversity. We hope that our survey, novel metric, and utility law will foster mutual advancement in both evaluation and mechanism interpretability [^1] Our code can be found at https://github.com/ALEX-nlp/MUI-Eval.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in both academic and industrial domains, largely attributed to the rapid advancements in pre-training and post-training techniques. While, current evaluation methods have still not kept pace with the progress in training — although various benchmarks and assessment for LLMs have emerged, few works focus on evaluation metric and mechanism, which leads to various problems in practices such as:
  • How can we select or curate a suitable dataset that covers the necessary capabilities of interest?
  • In what ways can the evaluation process effectively guide LLM training and refinement?
  • To what extent does data contamination undermine our evaluations?
  • Are current benchmarks fair in comparing different LLMs?
To systematically investigate the above issues, we revisit the core limitations of the conventional evaluation pipeline in the era of LLMs — the contradiction between the growing capabilities of models and limited evaluation data. As illustrated in Figure 1, by introducing more training data, compute, and parameters (i.e., scaling law), LLMs can generalize well across an increasing number of real-world tasks. While, considering the evaluation efficiency, we cannot expand the size of the testing dataset infinitely. This indicates a varied coverage of model abilities within the testing samples, resulting in unfair comparison among different models or different checkpoints of the same model. In specific, we define in-domain testing when samples fall within the model’s capacity, which are apparently easy to answer; conversely, out-of-domain testing suggests that the samples are beyond the model’s current capabilities, which are relatively difficult to answer. Thus, the more in-domain samples there are relative to out-of-domain ones, the better the model will appear to perform. That is, the more aligned the testing samples are with a model’s capabilities, the greater the advantage the model will have in the evaluation. Take the data contamination issue as an example, some models may experience overestimated performance just because the testing samples, which have become in-domain through pre-training or post-training, align perfectly with the model’s abilities. This, however, cannot reflect the model’s true capabilities. Therefore, how to quantitatively estimate the alignment between evaluation target and a model’s inherent capacities is a key challenge in gauging true model performance with limited data. We term this the evaluation generalization issue.
Figure 1: Illustration of how our metric MUI helps to mitigate the contradiction between the growing capabilities of models and limited evaluation data. 1) The right upper side: model capabilities increase, which naturally covers more in-domain testing and achieves higher performance, indicated by the same or lower MUI. 2) The right lower side: model capabilities remain unchanged or weakened, where higher performance is achieved by targeted optimization for more in-domain testing (e.g., data contamination), with higher utilization (MUI).
In this paper, we propose a new metric, Model Utilization Index (MUI), for generalizable evaluation by introducing the mechanism interpretability techniques. MUI calculates the proportion of neurons or features activated to complete given tasks, thereby enabling measurement of the degree to which a model utilizes its capabilities during evaluation. Combined with traditional performance metrics, we thus offers enhanced assessment comprehensiveness beyond limited testing data — this parallels human evaluation paradigms: when assessing an individual’s overall ability, we consider not only their task performance, but also the effort expended to achieve that outcome. Clearly, the less effort required to achieve a better result, the stronger the individual’s ability. In addition, MUI can also serve as an indicator of dataset diversity for a given model — optimal data variety should elicit maximal breadth of a model’s capacities, thereby preventing redundant activation of identical capacity subsets.
Based on MUI, we have conducted a series of empirical studies to demonstrate its effectiveness, which reveal a Model Utility Law and four corollaries of LLM evaluation. Specifically, given fixed datasets, MUI exhibits a negative logarithm relationship with model performance (Section 2). Building on it, our corollaries can address the questions raised earlier. The insights are summarized as follows: 1) When studying the training of a single model, an increase in MUI on a fixed dataset may signal potential degradation of unseen capabilities. This is useful for regression testing of the model’s quality using a limited set of testing samples. 2) When a model’s performance is overestimated due to data contamination, MUI does not decrease as it would in standard pre-training or post-training scenarios; rather, it tends to increase. 3) When comparing multiple models, stronger overall model capabilities should be indicated by both higher performance and lower MUI. This also explains why many models that perform similarly to GPT-4 on leaderboards often deliver vastly different user experiences: their higher MUI suggests they achieve similar performance by maximizing model utilization rather than improving fundamental capabilities. 4) Compared to traditional data diversity measurements like domains and capabilities, MUI is not only a comprehensive indicator, but it also emphasizes that diversity is model-specific. Of course, MUI is still limited by the current state of interpretability research. We look forward to the mutual advancement of these two fields.
We summarize our contributions as follows:
  • We highlight the generalization issue of current evaluation using limited data.
  • We propose a new metric MUI complementing traditional performance metrics.
  • We have concluded a utility law and four corollaries from extensive experiments, to shed light on future investigation in the fields of evaluation and interpretability.
@@ 3.2 Utility Law
Figure 3: Overall MUI-performance relationship across six datasets. MMLU is excluded due to DeepSeek series model inference cost considerations. According to model utilization curve, when performance reaches 100%, the minimum MUI is around 6.97%.
@@
where $\mathcal{P}$ denotes performance score, and we have $A=-4.632,B=28.305$ for overall relationship. We also observe similar trend using SAE-based MUI, which can be found in Appendix B.1.
From the above model utilization law, we can observe that: 1) From the upper left to the lower right, the ranking of model capabilities generally aligns with expectations. In particular, for models with similar performance, such as Vicuna and Llama2, MUI distinctly differentiates their ability levels. A more detailed analysis is provided in Section 3.5 (Corollary 3); 2) This theorem can also explain why MoE models (e.g., Deepseek R1) perform well, as they incorporate a sparsity objective during optimization, which is equivalent to minimizing MUI; 3) There are two special points on the logarithmic curve. First, as performance approaches zero, MUI tends toward infinity, implying that evaluation loses its meaning; second, when the performance score reaches 100%, MUI approaches approximately 6.97%. This may indicate a limit compression ratio — if a lower MUI is preferable, then this value represents the minimum expected model utilization, thus providing guidance for training. Of course, due to dataset limitations and the constraints of interpretability techniques, this value may not be precise. We will continue to explore this further.
@@ 3.3 MUI for Model Training
Regression Testing: During model training, an increase in model utilization on some dataset may indicate a decline in other capabilities beyond the dataset.
@@
Figure 5: MUI and Performance relationship, studied on Llama / Qwen series. We basically compare the math / code versions with the base models, to see the changes using in-domain testing and out-of-domain testing (e.g., for code version, MBPP is in-domain and GSM8K is out-of-domain). The orange arrow from the right lower side to the left upper side denotes the coarsening direction, and the blue arrow from the left lower side to right upper side denotes the accumulating direction.
We now empirically demonstrate these four directions based on existing LLMs, followed by investigating the pre-training progress of the OLMo series thanks to its open-source details. First, for the evolving direction, we have demonstrated it in Section 3.2. Similar trend can also be found in Figure 14, if we only include the same series of LLMs, e.g., vicuna, Llama2, Llama3.1, and Deepseek-distill-Llama. Second, for the coarsening and accumulating directions, we assume that these two typically occur during the process of enhancing specific abilities. We thus choose math reasoning and coding as two types of enhanced abilities, observing the Llama series: Llama-2-7B-Chat and CodeLlama-7B-Instruct, where the latter is tuned on coding data based on the former LLM, and the Qwen series: Qwen2.5-7B-Instruct and its coding/math optimized versions, Qwen2.5-Coder-7B-Instruct and Qwen2.5-Math-7B-Instruct. As shown in Figure 5, we can see that the specialized versions move in the accumulating direction compared with the base model if testing on the datasets for targeted abilities, e.g., CodeLlama vs Llama2 on MBPP, Qwen2.5-Coder vs Qwen2.5 on MBPP, and Qwen2.5-Math vs Qwen2.5 on GSM8k, while the specialized versions move in the coarsening direction compared with the base model if testing on the out-of-domain (OOD) benchmark, e.g., CodeLlama vs Llama2 on GSM8k, Qwen2.5-Coder vs Qwen2.5 on GSM8k, and Qwen2.5-Math vs Qwen2.5 on MBPP. We attribute the main reason is when the performance improvement on a particular task has not yet reached the level of fundamental capability enhancement—primarily due to an overall distributional shift towards that ability—it corresponds to the accumulating direction (with an increase in MUI and performance). Simultaneously, this shift results in decreased performance on other tasks, characterizing the coarsening direction (with an increase in MUI and a decrease in performance). Full results on the selected tasks shown in Table 5 and Table 9. This observation in turn suggests that using MUI as an indicator on one fixed dataset can relate to changes in other capabilities, which we summarize as Corollary 1. Third, for the collapsing direction, we only observe it when data contamination occurs, which will be detailed in Section 3.4.
@@
Figure 6: MUI-Performance curve for OLMo-2-7B trained using 0.5T, 1T, 2T, 3T 4T tokens on MATH and BBH.
@@ 3.5 MUI for Model Comparison
Vicuna-7B 11.9 / 5.1 4.0 / 1.7 54.0 / 19.6 13.4 / 12.2 23.6 / 17.6 28.6 / 9.9 9
Llama-2-7B-Chat 25.8 / 13.2 5.4 / 2.4 60.0 / 25.8 16.4 / 17.3 24.8 / 21.0 26.2 / 9.0 8
Qwen1.5-7B-Chat 58.7 / 33.9 17.9 / 9.4 76.2 / 30.6 39.0 / 41.1 39.8 / 33.6 41.3 / 14.2 7
Llama-3-8B-Instruct 79.5 / 43.8 26.6 / 11.7 82.1 / 34.4 59.8 / 54.6 56.5 / 41.0 64.0 / 23.9 6
Llama-3.1-8B-Instruct 86.5 / 52.6 48.1 / 24.7 81.9 / 42.0 62.2 / 56.8 57.9 / 39.0 65.4 / 23.6 5
Gemma-2-9B-Instruct 82.3 / 55.5 33.5 / 18.7 89.5 / 47.2 63.4 / 75.8 61.2 / 58.4 57.5 / 26.3 4
Qwen2.5-7B-Instruct 84.5 / 61.3 61.9 / 37.0 86.8 / 42.4 71.3 / 85.2 62.1 / 62.1 66.0 / 28.7 3
DeepSeek-Llama3.1-8B 71.6 / 40.7 74.1 / 39.1 82.6 / 43.5 68.9 / 72.6 60.3 / 49.2 76.8 / 33.4 2
DeepSeek-Qwen2.5-7B 82.6 / 51.2 80.1 / 42.2 81.1 / 44.0 71.9 / 85.9 62.5 / 57.1 66.5 / 30.4 1
Spearman 68.3 / 68.3 98.3 / 98.3 66.7 / 90.0 98.3 / 95.0 95.0 / 85.0 91.7 / 95.0 86.4Δ1.8 / 88.6Δ1.0
Kendall 55.6 / 61.1 94.4 / 94.4 50.0 / 83.3 83.3 / 94.4 88.9 / 72.2 77.8 / 83.3 76.9Δ3.2 / 80.6Δ1.2
@@
Our finds are as follows. 1) Higher correlation. The PUR-based ranking aligns more closely with the reference than the performance-based ranking, achieving average correlations of 88.6% versus 86.4% under Spearman, and 80.6% versus 76.9% under Kendall, thereby providing quantitative evidence of its effectiveness. 2) Lower variance. The PUR-based ranking exhibits smaller variance across datasets: 1.0 versus 1.8 (Spearman) and 1.2 versus 3.2 (Kendall), indicating that PUR delivers more stable ordering. 3) Qualitative improvements. Inspection of several ambiguous or previously mis-ranked cases (see Figure 3) shows how PUR resolves inconsistencies. For instance, although Vicuna and Llama-2 achieve nearly identical performance, MUI reveals Llama-2’s substantially greater underlying capability, allowing PUR to distinguish them; a similar clarification arises between Llama-3.1 and Qwen-2.5. Likewise, while Gemma-2 is generally regarded as stronger than Llama-3.1, a performance-only ranking suggests the opposite, whereas the PUR-based ranking corrects this by incorporating MUI.
@@ B.1 Utility Law: the Relationship Between MUI and Performance
Vicuna-7B 11.9 / 5.4 4.0 / 5.5 54.0 / 7.6 13.4 / 1.2 23.6 / 1.8 28.6 / 8.4 45.3 / 26.1
Llama-2-7B-Chat 25.8 / 3.8 5.4 / 4.9 60.0 / 5.4 16.4 / 0.9 24.8 / 1.4 26.2 / 8.5 48.2 / 15.8
CodeLlama-7B-Instruct 23.0 / 4.5 6.0 / 7.0 49.5 / 8.7 34.1 / 1.8 44.7 / 2.9 21.8 / 8.3 41.4 / 24.5
Qwen1.5-7B-Chat 58.7 / 3.0 17.9 / 3.6 76.2 / 6.2 39.0 / 0.9 39.8 / 1.4 41.3 / 8.5 58.3 / 19.8
Llama-3.1-8B-Instruct 86.5 / 2.7 48.1 / 3.8 81.9 / 3.8 62.2 / 1.2 57.9 / 2.2 65.4 / 7.7 64.4 / 14.5
Gemma2-9B-Instruct 82.3 / 2.2 33.5 / 3.2 89.5 / 3.6 63.4 / 0.7 61.2 / 1.1 57.5 / 4.8 73.3 / 10.2
Qwen2.5-7B-Instruct 84.5 / 1.9 61.9 / 2.8 86.8 / 4.2 71.3 / 0.7 62.1 / 1.0 66.0 / 5.3 71.3 / 14.5
Qwen2.5-Coder-7B-Instruct 79.6 / 3.7 49.5 / 6.2 83.8 / 9.4 75.0 / 1.5 65.9 / 2.4 57.8 / 11.0 65.6 / 30.3
Qwen2.5-Math-7B-Instruct 92.1 / 1.9 76.7 / 3.3 66.3 / 3.8 53.0 / 0.7 47.9 / 1.1 29.1 / 2.9 51.2 / 8.5
DeepSeek-Llama3.1-8B 71.6 / 3.1 74.1 / 3.6 82.6 / 3.6 68.9 / 0.9 60.3 / 1.5 76.8 / 5.3 -
DeepSeek-Qwen2.5-7B 82.6 / 2.6 80.1 / 3.6 81.1 / 3.4 71.9 / 0.7 62.5 / 1.2 66.5 / 4.8 -
Vicuna-13B 22.4 / 4.1 5.0 / 4.7 61.9 / 7.5 17.1 / 1.1 27.6 / 1.7 42.3 / 6.8 50.9 / 24.5
Llama-2-13B-Chat 36.2 / 3.5 7.4 / 4.1 66.8 / 5.0 18.9 / 0.9 29.7 / 1.4 26.1 / 7.9 53.6 / 14.0
Qwen1.5-14B-Chat 70.7 / 4.5 25.6 / 6.0 86.3 / 8.8 50.6 / 1.5 49.9 / 2.4 50.3 / 11.7 65.4 / 30.3
Qwen2.5-14B-Instruct 84.7 / 4.9 60.9 / 6.7 91.5 / 8.5 68.3 / 1.4 63.3 / 2.3 64.0 / 11.1 77.2 / 26.9
Qwen2.5-Coder-14B-Instruct 86.3 / 4.1 55.7 / 6.6 88.4 / 8.9 81.1 / 1.6 67.3 / 2.6 71.7 / 10.9 71.0 / 28.3
Vicuna-33B 35.7 / 3.1 8.1 / 4.3 67.1 / 9.4 17.7 / 1.1 25.1 / 1.6 46.0 / 6.6 51.5 / 34.8
Qwen2.5-32B-Instruct 86.7 / 2.9 67.3 / 2.1 92.2 / 5.6 72.0 / 1.0 69.9 / 1.4 76.4 / 3.7 78.5 / 19.1
@@
OLMo-2-7B-0.5T 7.6 / 6.5 3.8 / 5.7 52.3 / 6.5 4.3 / 0.8 8.2 / 1.2 23.3 / 6.4 42.5 / 17.9
OLMo-2-7B-1T 8.0 / 5.6 4.2 / 6.6 59.6 / 8.3 4.8 / 1.0 10.6 / 1.5 27.1 / 7.7 45.2 / 22.4
OLMo-2-7B-1.5T 10.8 / 6.4 4.7 / 8.5 62.2 / 9.2 9.7 / 1.1 14.8 / 1.7 27.3 / 8.9 49.1 / 24.7
OLMo-2-7B-2T 13.5 / 5.8 4.8 / 7.1 66.0 / 9.5 9.1 / 1.2 15.0 / 1.8 30.9 / 8.6 50.7 / 24.9
OLMo-2-7B-2.5T 15.2 / 6.6 5.5 / 7.1 66.0 / 9.5 6.7 / 1.2 15.8 / 1.9 35.0 / 9.2 51.2 / 24.5
OLMo-2-7B-3T 15.3 / 6.1 5.3 / 7.5 71.2 / 9.4 9.1 / 1.2 18.8 / 1.9 35.1 / 8.9 53.2 / 25.7
OLMo-2-7B-3.5T 16.8 / 5.4 5.7 / 6.5 71.4 / 8.8 8.5 / 1.1 20.8 / 1.7 38.5 / 8.2 53.7 / 23.8
OLMo-2-7B-4T (final) 68.2 / 4.4 15.4 / 6.3 79.8 / 9.0 19.5 / 1.1 29.3 / 1.7 46.5 / 8.2 59.2 / 23.4

### v3

# Model Utility Law: Evaluating LLMs beyond Performance through Mechanism Interpretable Metric

@@ Abstract
Large Language Models (LLMs) have become indispensable across academia, industry, and daily applications, yet current evaluation methods struggle to keep pace with their rapid development. One core challenge of evaluation in the large language model (LLM) era is the generalization issue: how to infer a model’s near-unbounded abilities from inevitably bounded benchmarks. We address this challenge by proposing Model Utilization Index (MUI), a mechanism interpretability enhanced metric that complements traditional performance scores. MUI quantifies the effort a model expends on a task, defined as the proportion of activated neurons or features during inference. Intuitively, a truly capable model should achieve higher performance with lower effort. Extensive experiments across popular LLMs reveal a consistent inverse logarithmic relationship between MUI and performance, which we formulate as the Utility Law. From this law we derive four practical corollaries that (i) guide training diagnostics, (ii) expose data contamination issue, (iii) enable fairer model comparisons, and (iv) design model-specific dataset diversity [^1] Our code can be found at https://github.com/ALEX-nlp/MUI-Eval.
@@ 1 Introduction
Recent years have witnessed an explosion of research interest in evaluating large language models (LLMs). Rigorous evaluation not only identify model weaknesses to guide optimization but also helps to define new objectives for next-generation models. Although many benchmarks have emerged, researchers have highlighted a generalization issue in evaluation [6]. With increasing data, compute, and parameters, LLMs’ capabilities can scale nearly without bound (i.e., scaling law), but the size of benchmarks remains bounded due to practical efficiency concerns. Consequently, a key challenge in the era of LLMs lies in estimating a model’s true performance and potential beyond what is explicitly reflected in limited testing samples.
In this paper, we propose a new metric, Model Utilization Index (MUI), for generalizable evaluation. The basic idea parallels human assessment: when judging an individual, we weigh not only what they achieve but also how much effort they expend. MUI targets the latter dimension for LLMs, complementary to conventional performance metrics. Clearly, the underlying assumption is that the less effort required to achieve a better result (i.e., lower MUI and higher performance score), the stronger the model’s true ability. However, how to quantify the model’s effort?
Inspired by recent advances in mechanistic interpretability, we define MUI as the measurement of the degree to which a model utilizes its capabilities during inference, calculated by the proportion of neurons or features activated to complete given tasks. In this field it is regarded that all of the model abilities, once it has been thoroughly trained, are encoded in neurons or in distinctive activation patterns across them. Neuron-level studies support this premise by showing that specific concepts are consistently localized [28]. When the model produces a given concept, the same subset of neurons contributes far more than the rest, responsible for utilizing that capability. Similarly, instead of localizing neurons, another line of works like sparse auto-encoders (SAEs) explains the model behaviors by mapping neuron activation patterns into a sparse feature space [20]. When the model expresses a particular skill, only a handful of features being activated. Thus, as illustrated in Figure 1, compared with a model’s overall capabilities, solving given tasks only utilizes a fraction of its neurons or features (the inner highlighted circle). Our MUI quantifies this ratio as the measurement of the model effort.
Figure 1: Illustration of model utility: how much effort (i.e., activated abilities) utilized to complete given tasks. Two example cases for generalizable evaluation: 1) A more capable model should achieve higher performance with less effort (lower MUI). 2) When data contamination occurs, higher performance is achieved by utilizing more effort (higher MUI), while model’s overall capabilities remain unchanged or even weakened.
By integrating MUI with performance metrics, we achieve a generalizable evaluation that transcends the limitations of bounded benchmarks, thereby alleviating various related assessment issues. We have conducted a series of experiments for verification. The insights are summarized as follows:
  • Model Utility Law. Agree with our assumption, MUI exhibits a negative logarithm relationship with performance. This also explains why sparse mixure-of-expert (MoE) model performs better – it is not merely because of larger-scale parameters, but the introduction of an additional optimization objective to lower model utility. Our law suggests a limit sparsity ratio (Section 3.2).
  • Corollary 1 for training diagnostics. An increase in MUI may signal potential degradation of unseen capabilities. We conclude four optimization directions based on MUI and performance curve for deeply analyzing the training process using a limited set of testing samples (Section 3.3).
  • Corollary 2 for data contamination analysis. Unlike standard pre-training or post-training, contamination inflates performance with increasing MUI, accompanied by a specific optimization trajectory (Section 3.4).
  • Corollary 3 for fair model comparison. More capable model should be indicated by both higher performance and lower MUI. This also explains why two models with similar leaderboard scores can differ markedly in user experience; higher MUI reveals those that rely on brute utilisation rather than enhanced fundamental capability (Section 3.5).
  • Corollary 4 for data diversity. For a fixed model, optimal data variety should elicit maximal breadth of a model’s capacities, a.k.a., higher MUI. Compared to traditional data diversity measurements like domains and capabilities, MUI is not only a comprehensive indicator, but it also emphasizes that diversity is model-specific (Section 3.6).
Of course, MUI is still limited by the current state of interpretability techniques, whose progress will continuously push forward a more precise model utilization evaluation.
@@ 3.2 Utility Law
Figure 3: Overall MUI-performance relationship across six datasets. MMLU is excluded due to DeepSeek series model inference cost considerations. According to model utilization curve, when performance reaches 100%, the minimum MUI is around 9.77%.
@@
where $\mathcal{P}$ denotes performance score, and we have $A=-3.534,B=26.049$ for overall relationship. We also observe similar trend using SAE-based MUI, which can be found in Appendix B.1.
From the above model utilization law, we can observe that: 1) From the upper left to the lower right, the ranking of model capabilities generally aligns with expectations. In particular, for models with similar performance, such as Vicuna and Llama2, MUI distinctly differentiates their ability levels. A more detailed analysis is provided in Section 3.5 (Corollary 3); 2) This theorem can also explain why MoE models (e.g., Deepseek R1) perform well, as they incorporate a sparsity objective during optimization, which is equivalent to minimizing MUI; 3) There are two special points on the logarithmic curve. First, as performance approaches zero, MUI tends toward infinity, implying that evaluation loses its meaning; second, when the performance score reaches 100%, MUI approaches approximately 9.77%. This may indicate a limit compression ratio — if a lower MUI is preferable, then this value represents the minimum expected model utilization, thus providing guidance for training. Of course, due to dataset limitations and the constraints of interpretability techniques, this value may not be precise. We will continue to explore this further.
@@ 3.3 MUI for Model Training
Training diagnostics: During model training, an increase in model utilization on some dataset may indicate a decline in other capabilities beyond the dataset.
@@
Figure 5: MUI and Performance relationship, studied on Llama / Qwen series. We basically compare the math / code versions with the base models, to see the changes using in-domain testing and out-of-domain testing (e.g., for code version, HumanEval is in-domain and GSM8K is out-of-domain). The orange arrow from the right lower side to the left upper side denotes the coarsening direction, and the blue arrow from the left lower side to right upper side denotes the accumulating direction.
We now empirically demonstrate these four directions based on existing LLMs, followed by investigating the pre-training progress of the OLMo series thanks to its open-source details. First, for the evolving direction, we have demonstrated it in Section 3.2. Similar trend can also be found in Figure 14, if we only include the same series of LLMs, e.g., vicuna, Llama2, Llama3.1, and Deepseek-distill-Llama. Second, for the coarsening and accumulating directions, we assume that these two typically occur during the process of enhancing specific abilities. We thus choose math reasoning and coding as two types of enhanced abilities, observing the Llama series: Llama-2-7B-Chat and CodeLlama-7B-Instruct, where the latter is tuned on coding data based on the former LLM, and the Qwen series: Qwen2.5-7B-Instruct and its coding/math optimized versions, Qwen2.5-Coder-7B-Instruct and Qwen2.5-Math-7B-Instruct. As shown in Figure 5, we can see that the specialized versions move in the accumulating direction compared with the base model if testing on the datasets for targeted abilities, e.g., CodeLlama vs Llama2 on HumanEval, Qwen2.5-Coder vs Qwen2.5 on HumanEval, and Qwen2.5-Math vs Qwen2.5 on GSM8k, while the specialized versions move in the coarsening direction compared with the base model if testing on the out-of-domain (OOD) benchmark, e.g., CodeLlama vs Llama2 on GSM8k, Qwen2.5-Coder vs Qwen2.5 on GSM8k, and Qwen2.5-Math vs Qwen2.5 on HumanEval. We attribute the main reason is when the performance improvement on a particular task has not yet reached the level of fundamental capability enhancement—primarily due to an overall distributional shift towards that ability—it corresponds to the accumulating direction (with an increase in MUI and performance). Simultaneously, this shift results in decreased performance on other tasks, characterizing the coarsening direction (with an increase in MUI and a decrease in performance). Full results on the selected tasks shown in Table 5 and Table 9. This observation in turn suggests that using MUI as an indicator on one fixed dataset can relate to changes in other capabilities, which we summarize as Corollary 1. Third, for the collapsing direction, we only observe it when data contamination occurs, which will be detailed in Section 3.4.
@@
Figure 6: MUI-Performance curve for OLMo-2-7B trained using 0.5T, 1T, 2T, 3T, 4T tokens on MATH and BBH.
@@ 3.5 MUI for Model Comparison
Vicuna-7B 11.9 / 4.9 4.0 / 1.6 54.0 / 18.5 13.4 / 11.8 23.6 / 16.7 28.6 / 9.6 9
Llama-2-7B-Chat 25.8 / 12.4 5.4 / 2.3 60.0 / 23.9 16.4 / 16.4 24.8 / 19.0 26.2 / 8.9 8
Qwen1.5-7B-Chat 58.7 / 30.5 17.9 / 8.3 76.2 / 28.2 39.0 / 34.2 39.8 / 28.9 41.3 / 13.5 7
Llama-3-8B-Instruct 79.5 / 39.8 26.6 / 10.5 82.1 / 30.4 59.8 / 48.8 56.5 / 36.5 64.0 / 21.8 6
Llama-3.1-8B-Instruct 86.5 / 46.2 48.1 / 21.5 81.9 / 36.6 62.2 / 50.8 57.9 / 35.2 65.4 / 21.8 5
Gemma-2-9B-Instruct 82.3 / 51.0 33.5 / 16.2 89.5 / 41.3 63.4 / 66.8 61.2 / 50.0 57.5 / 23.3 4
Qwen2.5-7B-Instruct 84.5 / 55.7 61.9 / 33.1 86.8 / 37.4 71.3 / 75.2 62.1 / 49.1 66.0 / 25.7 3
DeepSeek-Llama3.1-8B 71.6 / 36.3 74.1 / 33.8 82.6 / 36.9 68.9 / 65.7 60.3 / 43.7 76.8 / 28.4 2
DeepSeek-Qwen2.5-7B 82.6 / 48.5 80.1 / 38.6 81.1 / 39.6 71.9 / 71.9 62.5 / 49.4 66.5 / 26.7 1
Spearman 68.3 / 75.0 98.3 / 98.3 66.7 / 88.3 98.3 / 91.7 95.0 / 86.7 91.7 / 96.2 86.4Δ1.8  89.4Δ0.6
Kendall 55.6 / 66.7 94.4 / 94.4 50.0 / 77.8 94.4 / 83.3 88.9 / 72.2 77.8 / 87.3 76.9Δ3.2  80.3Δ0.9
@@
Our finds are as follows. 1) Higher correlation. The PUR-based ranking aligns more closely with the reference than the performance-based ranking, achieving average correlations of 89.4% versus 86.4% under Spearman, and 80.3% versus 76.9% under Kendall, thereby providing quantitative evidence of its effectiveness. 2) Lower variance. The PUR-based ranking exhibits smaller variance across datasets: 0.6 versus 1.8 (Spearman) and 0.9 versus 3.2 (Kendall), indicating that PUR delivers more stable ordering. 3) Qualitative improvements. Inspection of several ambiguous or previously mis-ranked cases (see Figure 3) shows how PUR resolves inconsistencies. For instance, although Vicuna and Llama-2 achieve nearly identical performance, MUI reveals Llama-2’s substantially greater underlying capability, allowing PUR to distinguish them; a similar clarification arises between Llama-3.1 and Qwen-2.5. Likewise, while Gemma-2 is generally regarded as stronger than Llama-3.1, a performance-only ranking suggests the opposite, whereas the PUR-based ranking corrects this by incorporating MUI.
@@ B.1 Utility Law: the Relationship Between MUI and Performance
Vicuna-7B 11.9 / 6.0 4.0 / 6.4 54.0 / 8.5 13.4 / 1.3 23.6 / 2.0 28.6 / 8.9 45.3 / 27.4
Llama-2-7B-Chat 25.8 / 4.3 5.4 / 5.5 60.0 / 6.3 16.4 / 1.0 24.8 / 1.7 26.2 / 8.7 48.2 / 17.6
CodeLlama-7B-Instruct 23.0 / 5.1 6.0 / 8.0 49.5 / 9.6 34.1 / 2.0 44.7 / 3.3 21.8 / 8.5 41.4 / 26.3
Qwen1.5-7B-Chat 58.7 / 3.7 17.9 / 4.6 76.2 / 7.3 39.0 / 1.3 39.8 / 1.9 41.3 / 9.4 58.3 / 21.3
Llama-3.1-8B-Instruct 86.5 / 3.5 48.1 / 5.0 81.9 / 5.0 62.2 / 1.5 57.9 / 2.7 65.4 / 9.0 64.4 / 18.0
Gemma2-9B-Instruct 82.3 / 2.6 33.5 / 4.3 89.5 / 4.7 63.4 / 0.9 61.2 / 1.5 57.5 / 6.1 73.3 / 14.6
Qwen2.5-7B-Instruct 84.5 / 2.3 61.9 / 3.5 86.8 / 5.4 71.3 / 0.9 62.1 / 1.6 66.0 / 6.6 71.3 / 17.7
Qwen2.5-Coder-7B-Instruct 79.6 / 4.3 49.5 / 7.5 83.8 / 11.1 75.0 / 1.8 65.9 / 3.0 57.8 / 13.1 65.6 / 34.4
Qwen2.5-Math-7B-Instruct 92.1 / 2.3 76.7 / 4.0 66.3 / 3.3 53.0 / 1.0 47.9 / 1.5 29.1 / 3.7 51.2 / 17.3
DeepSeek-Llama3.1-8B 71.6 / 3.9 74.1 / 4.8 82.6 / 5.0 68.9 / 1.1 60.3 / 1.9 76.8 / 7.3 -
DeepSeek-Qwen2.5-7B 82.6 / 2.9 80.1 / 4.3 81.1 / 4.2 71.9 / 1.0 62.5 / 1.6 66.5 / 6.2 -
Vicuna-13B 22.4 / 4.6 5.0 / 5.5 61.9 / 8.3 17.1 / 1.3 27.6 / 2.0 42.3 / 7.0 50.9 / 26.7
Llama-2-13B-Chat 36.2 / 4.1 7.4 / 4.9 66.8 / 6.0 18.9 / 1.1 29.7 / 1.6 26.1 / 8.2 53.6 / 16.5
Qwen1.5-14B-Chat 70.7 / 5.4 25.6 / 7.4 86.3 / 10.1 50.6 / 1.7 49.9 / 2.7 50.3 / 13.4 65.4 / 32.7
Qwen2.5-14B-Instruct 84.7 / 5.3 60.9 / 7.7 91.5 / 10.0 68.3 / 1.9 63.3 / 3.3 64.0 / 13.2 77.2 / 31.1
Qwen2.5-Coder-14B-Instruct 86.3 / 4.5 55.7 / 7.4 88.4 / 10.2 81.1 / 1.8 67.3 / 3.1 71.7 / 12.5 71.0 / 31.5
@@
OLMo-2-7B-0.5T 7.6 / 6.4 3.8 / 5.5 52.3 / 7.4 4.3 / 0.8 8.2 / 1.1 23.3 / 6.0 42.5 / 16.7
OLMo-2-7B-1T 8.0 / 4.3 4.2 / 6.5 59.6 / 8.2 4.8 / 1.0 10.6 / 1.5 27.1 / 7.2 45.2 / 20.7
OLMo-2-7B-1.5T 10.8 / 6.3 4.7 / 8.6 62.2 / 9.4 9.7 / 1.1 14.8 / 1.8 27.3 / 8.5 49.1 / 24.1
OLMo-2-7B-2T 13.5 / 4.8 4.8 / 7.3 66.0 / 9.4 9.1 / 1.2 15.0 / 1.9 30.9 / 8.3 50.7 / 23.9
OLMo-2-7B-2.5T 15.2 / 6.8 5.5 / 7.3 66.0 / 9.8 6.7 / 1.2 15.8 / 2.0 35.0 / 8.9 51.2 / 24.3
OLMo-2-7B-3T 15.3 / 6.1 5.3 / 7.6 71.2 / 9.6 9.1 / 1.3 18.8 / 2.0 35.1 / 8.8 53.2 / 25.6
OLMo-2-7B-3.5T 16.8 / 5.4 5.7 / 6.5 71.4 / 9.1 8.5 / 1.2 20.8 / 1.9 38.5 / 8.0 53.7 / 23.8
OLMo-2-7B-4T (final) 68.2 / 5.1 15.4 / 6.5 79.8 / 9.4 19.5 / 1.2 29.3 / 1.8 46.5 / 8.2 59.2 / 24.0